# Functional Eigen-Grasping Using Approach Heatmaps

### v1

@@ I Introduction
The key idea of realizing the functional grasp is to utilize the heatmap of the palm position combined with the eigengrasp [6]. Setting one of the fingers as the functional finger to realize a function associated with the object, we obtain the heatmap of the palm position on the object surface, called the approach heatmap. The approach heatmap describes ideal locations on the object for the robotic ”palm” to grasp from that ensures the ability of the functional finger to contact the functional part of the object and perform the function. This approach heatmap is obtained using the directional manipulability of the functional finger. By referring to the approach heatmap, we can easily obtain the position of the palm to realize the function associated with the object. Once the position of the palm can be determined, the remaining fingers are used just to grasp the object firmly. For planning the grasping pose of the remaining fingers with low-dimensional control inputs, we utilize the eigengrasp. Here, different from the conventional eigengrasp, we apply three kinds of energy functions, i.e., the palm energy making the palm contact the object surface, the contact energy making the remaining fingers contact the object surface, and the functional energy making the functional finger contact the object surface.
@@ II Related Works
On the other hand, functional grasping has only been proposed by a few researchers. By focusing on human-to-robot mapping, Handa et al. [1] proposed a teleoperation method for controlling the robotic grippers, [2] [16] uses learning from demonstration, teaching the robot how to perform a particular task from a few examples of the human performing the same task. Nonetheless, these grasp retargeting approaches rely on handcrafted correspondences between the human hand and the robotic gripper, such as joint mapping, fingertip pose retargeting, and hand posture recognition-based mapping.
ContactGrasp [3] generates functional grasp by employing a sample and rank approach, wherein numerous grasps are sampled on an object and filtered based on their resemblance derived from human demonstrations [17]. Du et al. [5] proposed to align the robotic gripper with the human grasp’s position and orientation. Mandikal [18] built a model to learn grasp policies that predict the object’s grasp affordance based on contact heatmap data. Zhu et al. [19] proposed to annotate functional parts of the object for grasp planning. In all the above methods, the functional grasp has been generated by imitating human grasps. However, mapping joints and links is not trivial for application to non-anthropomorphic hands.
@@
III Assumption
A multi-fingered hand is tasked with achieving functional grasping of a tool, which includes specific functional elements such as triggers and buttons. The initial step involved manually selecting one of the fingers as the functional finger, which is responsible for manipulating the functional part of the tool. We assume that the functional finger has at least 2 DOFs, while the remaining fingers are assumed to have DOFs capable of stably grasping the tool using the eigengrasp. Additionally, we assume that each finger and palm makes soft-finger contact with the tool.
Subsequently, we introduce a constraint whereby the distal part of the functional finger makes contact with the functional part of the object. Referred to as the ’functional constraint,’ this constraint serves as a guiding principle for the robotic gripper during its object exploration. As illustrated in Fig. 2, the functional constraint serves a dual purpose, contributing to both creating the approach heatmap and informing the grasp energy formulation. This additional constraint is incorporated alongside conventional grasping metrics to ensure the functionality of the resulting grasp.
IV Method Overview
Our methodology for generating functional grasps runs in two phases. In the offline phase, we create an approach heatmap by first sampling the object for reachability from the functional meeting point - the intersection between the functional part of the gripper and the object. The inverse kinematic solution is then executed to determine reachability, and each point undergoes grading based on the directional manipulability index, with a subsequent refinement using a clustering algorithm.
During runtime, the grasp planning process utilizes eigengrasp. The simulated annealing planner optimizes three energy functions: the palm energy functional from the approach heatmap, the functional energy term ensuring alignment between the functional part of the object and the gripper, and the contact energy term for secure holding of the remaining portions. This intricate interplay results in a functional grasp. Fig. 2 provides a visual representation of our method.
V Constructing the Approach Heatmap
@@
V-A Sampling Functional Reachability
@@
To establish the grasp affordance regions, the initial step involves identifying which regions on the object are reachable by the palm while adhering to the functional constraint. This process is carried out within a simulated environment, where multiple potential approach points are sampled on the object’s surface. We assess the feasibility of each of these points for palm access while remaining within the confines of the functional constraint.
Addressing inverse kinematics (IK) problems for dexterous robotic grippers can be notably complex due to the many joints involved. However, when focusing on the reachability of the functional finger to the functional part of the object, our concern primarily revolves around palm movement to the functional part of the object. Following this rationale, the IK problem is simplified to a chain of links from the tip of the functional finger to the robotic palm.
@@ V-B Directional Manipulability
Furthermore, we use the weighted least-norm method mentioned [24] to penalize configurations close to the joint limit. We focus on the linear component of the resulting Jacobian, extracting the eigenvalues and eigenvectors.
@@
This approach enables a comparative analysis of how effectively the robot can interact with each object at different joint configurations. By specifying the direction of the function t, we use the equation to grade each point at each joint configuration by how likely they are to move in that direction.
@@
V-C Refining the Approach Heatmap
The heatmap is further refined by applying the HDBSCAN clustering algorithm [25], removing any noise present in the heatmap. This improves performance, as multiple peaks in the gradient calculations can adversely impact the efficiency of the grasping algorithm. We opt for the largest cluster with the highest score on the object. This ensures that the selected cluster not only represents a substantial portion of the heatmap but also possesses the highest potential for functional grasping, thereby further optimizing the method’s overall performance. Fig. 1 gives a visual representation of this refinement, where the original heatmap is depicted in blue on the object, while the clustered region is highlighted in red.
VI Functional Grasp Planning
@@
VI-A Contact and Functional Energy
We have implemented the contact energy function introduced in [6]. This energy function assesses the proximity of the robotic hand to the object being grasped. As illustrated in Fig. 5, by defining virtual contact points on the robotic gripper, it can quantify the alignment between each of these points and the nearest point on the object, considering both the distance and the dot product of the normalized vector extending from the virtual contact location to the closest point on the object and the normal vector of the virtual contact point on the hand. Thus penalizing grasps distant from the object or where the normals are poorly aligned.
@@ VI-A Contact and Functional Energy
On the other hand, the functional energy function incorporates the functional constraint into the grasp framework. It operates similarly to the energy function described above but focuses exclusively on the virtual contact points located on the functional part of the robot, assessing their alignment with the functional area of the object. In this case, lower energy function values also indicate better performance.
VI-B Palm Energy
@@
6:     if $score>\alpha$ then
@@
This segment of the energy function is derived from the generated approach heatmap, establishing a geometric constraint that specifies a particular region of the object to be grasped. As outlined in Algorithm 2, this energy function calculates the distance between the robot’s palm and the nearest area on the object while considering the area’s score from the approach heatmap. If the calculated value falls below a predetermined threshold, it returns the maximum energy value, indicating that the current object region is unsuitable for functional grasping. Otherwise, it returns the average score of that region on the object from the approach heatmap, which is further adjusted by a factor, gamma. This multiplication penalizes moving away from regions with good scores. A higher gamma value encourages the palm pose to remain in areas with favorable scores, and the distance component ensures the palm remains near the object. This energy function is biased towards palm poses close to the higher-score regions of the object in the approach heatmap, resulting in lower energy configurations the closer they are to these areas.
VI-C Hybrid Energy
@@
Figure 7: Real World Experiments: Showcasing demonstrations of the functional grasps in real life using the Shadow Hand Lite attached to the Motoman SDA5F arm
VII Experimental Results
VII-A Approach Heatmap Generation
Approach heatmaps are generated for the anthropomorphic Shadow Hand Lite and the non-anthropomorphic Barrett Hand across the entire object set. The choice of the functional finger for the Shadow Hand Lite is intuitive, resembling the human hand, with the index finger selected in most cases, except for the remote control, where the thumb is chosen. However, for the Barrett Hand, the choice is less intuitive. Considering its Degrees of Freedom, we opt for one of the coupled fingers, mimicking the location of the index finger in the human hand.
A diverse selection of everyday objects is employed for testing, including varied designs for some objects. Einscan Pro 2x Plus is used to capture the geometrical model of these objects. Subsequently, each object is sampled into 500 potential approach points and evaluated based on weighted directional manipulability. Following the sampling, all scores undergo normalization by division with the 99th quantile score. This one-time offline procedure for each object requires approximately 15 minutes using a PC with Intel core i7-9750H @ 2.60 GHz CPU.
@@ VII-A Approach Heatmap Generation
VII-B Planning Functional Grasps
Using the generated approach heatmaps, the simulated annealing planner of eigengrasp was run for 70,000 steps using the Hybrid energy function, with the lowest energy grasp selected. For the Shadow Hand Lite, the energy function weights were allocated as $\alpha=0.75$, $\beta=0.2,$ and $\gamma=0.05$, while for the Barrett Hand, the weights were $\alpha=0.6$, $\beta=0.35$, and $\gamma=0.05$. The smaller contact energy weight for the Barrett Hand was due to its finger size, posing challenges in fully enveloping the object while attempting to fulfill the intended function.
Fig. 6 shows the experimental results of planning functional grasps, showing functional grasps for each object using the approach heatmaps. In each robotic gripper column, we present the resulting grasp alongside the approach heatmap utilized in that grasping process. Success was determined if the robotic gripper could make contact with the functional finger while maintaining force closure on the rest of the object with the remaining fingers. The functional energy term ensures alignment with the functional parts of the object, while force closure is assessed using the Grasp $\epsilon$-quality and $\nu$-quality metric. Table I exemplifies stability results for some grasps in Fig. 6. To assess the consistency of the generated grasps using the energy functions, the grasp planner underwent five iterations for each approach heatmap. Table II summarizes the overall consistency rate for each robotic gripper.
TABLE I: Stability of the grasps generated
Shadow Hand Lite Barrett Hand
object
$\epsilon$
$\nu$
$\epsilon$
$\nu$
a
0.179
2.23
0.0663
0.736
d
0.0975
2.74
0.149
1.74
f
0.286
21.6
0.353
25.9
TABLE II: Success rate of achieving the functional grasp in GraspIt! using the approach heatmap
Robotic Gripper Success Rate
Shadow Hand Lite 0.85
Barrett Hand 0.77
@@ VII-B Planning Functional Grasps
VII-C Real World Experiments
This study focuses on generating functional grasps, detailing a robotic gripper’s pose and joint configuration to enable it to perform an object’s function. Robotic perception falls beyond the scope of this research and can be addressed using alternative methods [26] [27] for hand and object pose estimation. To emphasize the practicality of our approach, we simplify the process and illustrate its feasibility. Our experiments employ the Shadow Hand Lite gripper mounted on the Motoman SDA5F arm, targeting objects such as a spray bottle, nozzle spray, and drill. In these experiments, the pose of the target objects is predetermined and provided to the robot system. The objective is to validate the physical feasibility of the grasps generated by our method.
Fig. 7 shows the executed grasps on the objects. All the objects, except the remote control, are placed on the table with their poses known to the robot. The robotic arm approaches the object to the target pose and gradually closes in on it. Using the tactile sensors on the fingertips, the functional finger lightly taps on the functional part of the object while the remaining fingers securely hold the object. In the case of the remote control, requiring a grasp involving picking up and in-hand manipulation beyond the paper’s scope, we manually hand over the remote to the robotic gripper while it closes.
VIII Conclusion
This paper presents a method for achieving functional grasps across diverse objects with different grippers. By defining the functional components of both the robotic gripper and the object, we propose the integration of a functional constraint to narrow down the grasp planning search space, focusing solely on grasps that fulfill the intended function. Additionally, we introduce the concept of the approach heatmap, revealing optimal locations on the object for the robotic palm to approach, enabling the gripper to satisfy the function effectively. Using the approach heatmap, we show experiments of planning and executing functional grasps on everyday objects.
This method enables each robotic gripper to identify its own suitable grasping areas, accommodating variations in kinematic design and capabilities. Currently, the generation of the approach heatmap and the grasp planning process are separate from each other. The grasp planner utilizes the approach heatmap in planning, influencing the types of objects the robotic gripper can grasp. However, this adherence to typical grasp planner metrics may have negative implications, particularly for objects like scissors that can’t be effectively grasped using the entire gripper. For future works, we will consider combining the two steps, enabling the robotic gripper to optimize based on both factors simultaneously. Also, we will consider the motion that comes with executing the function. We believe considering these factors would contribute to the generalization and performance of this method.

### v2

@@ I Introduction
The key idea of realizing the functional grasp is to use the palm position heatmap combined with the eigengrasp [6]. By designating a functional finger to interact with the object, we obtain the approach heatmap, which indicates ideal palm positions on the object’s surface. This ensures the functional finger can contact the functional part of the object. The approach heatmap is derived from the directional manipulability of the functional finger.
By referring to this heatmap, we can easily determine the palm’s position to achieve the object’s associated function. Once the palm position is set, the remaining fingers firmly grasp the object. For planning the grasping pose of these fingers with low-dimensional control inputs, we use eigengrasp, applying three energy functions: palm energy for palm-object contact, contact energy for the remaining fingers, and functional energy for the functional finger.
@@ II Related Works
On the other hand, functional grasping has only been proposed by a few researchers. By focusing on human-to-robot mapping, Handa et al. [1] proposed a teleoperation method for controlling the robotic grippers, [2] [16] uses learning from demonstration, teaching the robot how to perform a particular task from a few examples of the human performing the same task. Ye et al. [17] leverage implicit functions and Conditional Variational Autoencoders trained on retargeted human demonstration data. Nonetheless, these grasp retargeting approaches rely on handcrafted correspondences between the human hand and the robotic gripper, such as joint mapping, fingertip pose retargeting, and hand posture recognition-based mapping.
ContactGrasp [3] generates functional grasp by employing a sample and rank approach, wherein numerous grasps are sampled on an object and filtered based on their resemblance derived from human demonstrations [18]. Du et al. [5] proposed to align the robotic gripper with the human grasp’s position and orientation. Mandikal [19] built a model to learn grasp policies that predict the object’s grasp affordance based on contact heatmap data. Zhang et al. [20] proposed a semantic representation of functional hand-object interaction. They could label grasps by annotating specific parts of the hands, guiding their framework to generate functional grasps. Agarwal et al. [21] used internet data to learn a category-level policy for grasping objects based on their affordance areas. In all the above methods, the functional grasps have been generated by imitating human grasps. However, mapping joints and links is not trivial for application to non-anthropomorphic hands.
@@
III Method Overview
In this methodology, we introduce a ”functional constraint,” ensuring the distal part of the functional finger contacts the functional part of the object. This constraint guides the robotic gripper during object exploration. Our approach to generating functional grasps involves two phases.
In the offline phase, we create an approach heatmap by sampling the object’s reachability from the functional meeting point—the intersection of the gripper’s functional part and the object. We then solve the inverse kinematics to determine reachability, grading each point based on the directional manipulability index and refining them with a clustering algorithm.
During runtime, the grasp planning process employs eigengrasp. The simulated annealing planner optimizes three energy functions: the palm energy from the approach heatmap, the functional energy for alignment between the gripper and the object, and the contact energy for secure holding of the object. This intricate interplay results in a functional grasp. Figure 2 provides a visual representation of our method.
IV Constructing the Approach Heatmap
@@
IV-A Sampling Functional Reachability
@@
To establish the grasp affordance regions, we first identify which object regions the palm can reach while meeting the functional constraint. This is done in a simulated environment by sampling multiple potential approach points on the object’s surface and assessing their feasibility for palm access within the functional constraint.
Solving inverse kinematics (IK) for dexterous robotic grippers is complex due to the many joints involved. However, focusing on the reachability of the functional finger to the functional part of the object simplifies the IK problem to a chain of links from the tip of the functional finger to the robotic palm.
IV-B Directional Manipulability
@@
Furthermore, we use the weighted least-norm method mentioned [26] to penalize configurations close to the joint limit. Our analysis centers on the linear component of the resulting Jacobian matrix, from which we extract its eigenvalues and eigenvectors. The number of eigenvalues, denoted as $m$, corresponds to the dimensions of the Jacobian matrix, which in turn depends on the degrees of freedom of the robot’s functional link.
@@
This approach enables a comparative analysis of the robot’s effectiveness in interacting with each object at different joint configurations. By specifying the function direction $t$, which is defined along with the functional part (e.g., the rotational joint of a trigger or the direction of button pressing), we can use the equation to grade each point at each joint configuration based on its likelihood to move in that direction, as shown in Fig. 4.
@@
IV-C Refining the Approach Heatmap
The heatmap is refined using the HDBSCAN clustering algorithm [27] to remove noise, improving performance by preventing multiple peaks from affecting the grasping algorithm. We select the largest cluster with the highest score on the object, ensuring it represents a significant portion of the heatmap and has a high potential for functional grasping, optimizing overall performance. Fig. 1 shows this refinement, with the original heatmap in blue and the clustered region in red.
V Functional Grasp Planning
@@
V-A Contact and Functional Energy
We implemented the contact energy function from [6], which evaluates the robotic hand’s proximity to the object. As shown in Fig. 5, virtual contact points on the gripper measure alignment with the nearest object points by considering both distance and the dot product of vectors between them and the hand’s contact normals. This penalizes grasps that are distant or poorly aligned.
Contact energy is preferred over force closure metrics for achieving functional grasps. While force closure metrics overly focus on the object’s center of gravity, leading to suboptimal grasps, contact energy offers a more effective method for reliable and functional grasps.
@@ V-A Contact and Functional Energy
Additionally, the functional energy function incorporates the functional constraint into the grasp framework. It operates similarly to the energy function described above but focuses exclusively on the virtual contact points located on the functional part of the robot, assessing their alignment with the functional area of the object. In this case, lower energy function values also indicate better performance.
V-B Palm Energy
@@
6:     if $score>\omega$ then
@@
This energy function segment, derived from the approach heatmap, establishes a geometric constraint specifying where to grasp the object. As outlined in Algorithm 2, it calculates the distance between the robot’s palm and the nearest area on the object, considering the area’s heatmap score. If the score falls below a threshold $\omega$, it returns the maximum energy value, indicating the region is unsuitable for grasping. Otherwise, it returns the area’s average score, adjusted by a factor, gamma, which penalizes moving away from high-score regions. A higher gamma keeps the palm in favorable areas, ensuring proximity to the object and favoring palm poses near high-score regions for lower energy configurations.
V-C Hybrid Energy
@@
VI Experimental Results
VI-A Approach Heatmap Generation
Approach heatmaps are generated for the anthropomorphic Shadow Hand Lite and the non-anthropomorphic Barrett Hand across the entire object set. The functional finger for the Shadow Hand Lite is chosen intuitively to resemble the human hand. The index finger is used for some objects, as shown in Fig. 6, while the thumb is selected for others, such as the remote control in Fig. 9. For the Barrett Hand, the choice is less intuitive. Considering its Degrees of Freedom, one of the coupled fingers is chosen, mimicking the location of the index finger in the human hand.
We did not use a specific object dataset in our experimentation because existing datasets typically lack a variety of functional objects. Instead, we chose a group of everyday objects with multiple designs for testing and generated 3D models of these objects using the Einscan Pro 2X Plus scanner.
Each object was then sampled into 500 potential approach points and evaluated based on weighted directional manipulability. All scores were normalized by dividing them by the 99th quantile score. This one-time offline procedure for each object took approximately 15 minutes on a PC with an Intel Core i7-9750H @ 2.60 GHz CPU.
@@ VI-A Approach Heatmap Generation
Figure 7: Evaluation Metric: Assess the generated grasps by having the robotic gripper approach and grasp each object using the grasping fingers for stability, then lift the object and use the functional finger to interact with its functional part.
Figure 8: Real World Experiments: Demonstrating functional grasp experiments in real life using the Shadow Hand Lite attached to the Motoman SDA5F robot.
VI-B Planning Functional Grasps
Using the generated approach heatmaps, the simulated annealing planner of eigengrasp was run for 70,000 steps using the Hybrid energy function, with the lowest energy grasp selected. For the Shadow Hand Lite, the energy function weights were allocated as $\alpha=0.75$, $\beta=0.2,$ and $\gamma=0.05$, while for the Barrett Hand, the weights were $\alpha=0.6$, $\beta=0.35$, and $\gamma=0.05$.
TABLE I: Consistency Rates in GraspIt!
Object Shadow Hand Lite Barrett Hand
a 100% 100%
b 40% 100%
c 100% 80%
d 60% 80%
e 100% 60%
f 80% 100%
g 100% -
The differing values of the weights for the Shadow Hand and the Barrett Hand arise from their distinct designs. The Barrett Hand, with its smaller finger size, has a reduced contact weight, making it more challenging to fully envelop the object while still performing the intended function.
Figure 9: In real-world demonstrations, the thumb acts as the functional finger. For tasks like using a remote control, requiring extensive pre-grasp manipulation, the object is handed to the robotic gripper for a functional grasp.
Fig. 6 presents the experimental results of planning functional grasps, highlighting the functional grasps for each object with corresponding approach heatmaps. Each column represents a robotic gripper and displays the resulting grasp alongside the approach heatmap used during the grasping process. Additionally, we illustrate the use of the thumb as a functional finger in grasping a remote control in Fig. 9, demonstrating the versatility and effectiveness of our approach.
To evaluate the consistency of generating functional grasps using approach heatmaps in GraspIt! and to assess the energy functions, we had the grasp planner perform five attempts to generate a functional grasp for each object using the approach heatmap. We then recorded the number of successful grasps from these five attempts.
Table I summarizes the consistency of each robotic gripper across all objects. Overall, most objects maintained a very high success rate. However, some objects had a lower success rate due to a combination of factors, including object shape, the accessibility of the functional part, and the robot synergy’s ability to cover the object effectively. This suggests that running the grasp planner multiple times may be necessary to achieve successful results for these particular objects.
@@ VI-B Planning Functional Grasps
VI-C Real World Evaluation
Following the planning experiments, we evaluated our grasping framework using the Shadow Hand Lite mounted on a Motoman SDA5F robotic arm. The primary objective was to assess the framework’s capability to generate stable and functional grasps in a real-world setting.
Success was determined by the gripper’s ability to lift the object steadily, using non-functional fingers for grasping while pressing the designated functional part with the functional finger. Figure 7 demonstrates the gripper successfully pressing the functional button, provided the button’s stiffness is manageable.
The Shadow Hand Lite, constrained by its limited force generation, was unable to exert enough force to press the button due to its stiffness. This limitation is discussed in [28], suggesting that a robotic gripper with greater force capabilities could potentially resolve this issue.
As illustrated in Fig. 7, the gripper approaches the object based on its pose, grasps it with the non-functional fingers, lifts it, and then presses the functional finger onto the functional part. Our results demonstrate the effectiveness of this approach, with the robotic system successfully lifting all tested objects, as shown in Fig. 8.
VII Conclusion
This paper presents a method for achieving functional grasps across diverse objects using different grippers. By defining the functional components of both the robotic gripper and the object, we integrate a functional constraint to narrow the grasp planning search space, focusing solely on grasps that fulfill the intended function. We introduce the concept of the approach heatmap, which identifies optimal locations on the object for the robotic palm to approach, enabling effective functional grasps. Our experiments demonstrate planning and executing functional grasps on everyday objects using this heatmap.
This method allows each robotic gripper to identify suitable grasping areas, accommodating variations in kinematic design and capabilities. Currently, the approach heatmap generation requires the complete model and specification of the object’s functional part. To generalize this process to novel and previously unseen objects, we aim to automate it using methods like open-vocabulary part segmentation [29], which can detect the functional part from an RGB image. A key challenge will be integrating the concept of functional reachability to generate the heatmap.
Another area for improvement is in integrating the approach heatmap generation and grasp planning processes, as their separation limits the grasp planner’s effectiveness, especially for objects like scissors. Future work will focus on integrating these steps to optimize performance and generalize our method, considering the execution of functional motions. We believe this will enhance the generalization and performance of our method.